people to either Speech A or a different speech that is known not to impact the number of racist remarks people make. After each group hears their speech, Emily follows people in the groups and records the number of racist remarks made. Emily follows people not knowing which speech, they heard. a. What type of trial design is this? (1 point) b. Is this study blinded? (1 point) Please circle one: Yes No We do not know Ronin is studying the association between drug A and Bird Flu. He wants to determine if Drug A is superior to Drug B. He sees 100 patients in the first two days, and prescribes all of them Drug A. Then in the next three days, he prescribes 100 patients to Drug B, the standard of care. After one year for each group, 23 of those who took Drug B did not get bird flu and 40 of those who took Drug A did get bird flu. a) What is the alternative hypothesis for this study? (1 point) b) What is the cumulative incidence ratio of bird flu across both Drug A and B takers? (1 point) c) How might randomization have helped to reduce confounding in this example? (2 points)
Answer #1

Question a:

It is cross sectional design. Because it compare the different population at single point of time.

Question b:

No, this trial not blinded.

II.

Question a:

Drug B is having effectiveness in controlling bird flu.

Question b:

Cumulative incident= Number of cases / total numbers at risk

For drugs A= 23/100 = 0.23

For dry B= 40/ 100 = 0.4

Question 4:

Randomization reduce the entry of individuals with confounding Factors( risk bias in itself).